# Fading ergodicity

### v1

@@ Acknowledgements.
Figure S3: PDFs $P(|O_{nn}|)$ of the diagonal matrix elements of observables: (a) $\hat{O}=\hat{S}^{z}_{L}$ and (b) $\hat{O}=\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$, for the system size $L=12$. Dashed lines indicate fits to the Gaussian PDF from Eq. (S3). We numerically subtract the smooth structure function $O(\bar{E})$ to study the distribution of fluctuations.
Figure S4: Measures of gaussianity for the diagonal matrix elements of $\hat{S}_{L}^{z}$ vs $\alpha$, for different system sizes $L$. (a) The normalized variance $\Gamma_{O}-\pi/2$ from Eq. (S4), (b) the Binder cumulant $B-1$ from Eq. (S5), and (c) the Kurtosis $K-3$ from Eq. (S6). Horizontal dotted lines are the results for the Gaussian PDF. Vertical dashed lines denote the transition point $\alpha_{c}$ obtained from Fig. 3(d) in the main text.
Figure S5: PDFs of off-diagonal matrix elements of the observable $\hat{O}=\hat{S}_{L}^{z}$, studied in the energy interval $\omega<100\Delta$, for (a) $\alpha=0.86$ and (b) $\alpha=0.92$. Main panels: $P(|O_{nm}|)$ vs $|O_{nm}|$. Dashed lines are the fits of the Gaussian PDF from Eq. (S3) to the largest system size. Insets: $P(\ln|O_{nm}|)$ vs $\ln|O_{nm}|$. Solid lines are the fits of the corresponding Gaussian PDF from Eq. (S7) using the same variance.
Figure S3 shows the distributions of the absolute values of diagonal matrix elements of observables $\hat{S}_{L}^{z}$ and $\hat{S}_{\rm dot}^{z}\hat{S}^{z}_{L}$, for a fixed system size and for different values of $\alpha$. Their main feature is that the distribution broadens upon decreasing $\alpha$, which is consistent with the softening of fluctuations in the regime of fading ergodicity.
The dashed lines in Fig. S3 represent the Gaussian probability density functions (PDFs). Given a positive random variable $z$, the corresponding Gaussian PDF is given by
$\bar{P}(z)=\frac{2}{\sqrt{2\pi\sigma^{2}}}\exp{-\frac{z^{2}}{2\sigma^{2}}}\;,$ (S3)
@@
While the Gaussian PDFs in Fig. S3 provide accurate descriptions for the numerical distributions at large $\alpha$, they also exhibit deviations at lower $\alpha$ in the tails of the distributions. To quantify the deviations from the Gaussian PDF as functions of both $\alpha$ and $L$, we compute the normalized variance leblond_mallayya_19
$\Gamma_{O}=\frac{\expectationvalue{z^{2}}}{\expectationvalue{z}^{2}}\;,$ (S4)
which yields $\Gamma_{O}=\pi/2$ for a Gaussian PDF from Eq. (S3). We complement this measure by computing the Binder cumulant
$B=\frac{\expectationvalue{z^{4}}}{3\expectationvalue{z^{2}}^{2}}\;,$ (S5)
@@
$K=\frac{\expectationvalue{(z-\expectationvalue{z})^{4}}}{\expectationvalue{(z-\expectationvalue{z})^{2}}^{2}}\;,$ (S6)
@@
Figure S4 shows the dependence of the gaussianity measures from Eqs. (S4)-(S6) on $\alpha$, for the observable $\hat{S}_{L}^{z}$ and different system sizes $L$. We observe that the PDFs are indeed close to a Gaussian PDF at large $\alpha$, and notable deviations occur when the transition point is approached. However, the dependence on system size shares certain similarities with other ergodicity measures such as the average gap ratio $r$ studied in Fig. S1(b). This suggests a possibility that in the thermodynamic limit $L\to\infty$, the PDFs are close to Gaussian in the entire ergodic phase. The vertical dashed line in Fig. S4 represents the critical value $\alpha_{c}$ obtained by fitting the divergence of the fluctuation exponent $\eta$ in Fig. 3 using Eq. (10). Notably, all the gaussianity measures studied in Fig. S4 exhibit a scale invariant point that is very close to the extracted value of $\alpha_{c}$ from Fig. 3.
@@
$\tilde{P}(x)=\frac{2e^{x}}{\sqrt{2\pi\sigma^{2}}}\exp{-\frac{e^{2x}}{2\sigma^{2}}}\;.$ ) . (S7)
The results from the insets of Fig. S5 reinforce the observation from the main panels that the deviations from the Gaussian PDF occur mostly in the tails, i.e., for large values of matrix elements. The deviations are larger in Fig. S5(a) [at $\alpha=0.86$] than in Fig. S5(b) [at $\alpha=0.92$], suggesting similar finite-size phenomenology as observed in Fig. S4 for the diagonal matrix elements.
S5 Quantum dynamics of observables
We now focus on ergodicity through the lens of quantum dynamics. Our goal is to study the dynamics of the expectation value of observable $\hat{S}_{L}^{z}$ after a quantum quench,
$\mathcal{Q}(t)=\langle\psi_{0}|\hat{S}_{L}^{z}(t)|\psi_{0}\rangle\;.$ (S8)
We choose the initial state $|\psi_{0}\rangle$ such that we minimize the finite-size effects due to the existence of a many-body mobility edge in the quantum sun model pawlik_sierant_2024. Specifically, for a given Hamiltonian realization, denoted as $\hat{H}^{(\mu)}$, we select a single initial state $|\psi_{0}\rangle$ that is a product state in the computational basis $|i\rangle$ (i.e., an eigenstate of $\hat{S}_{L}^{z}$), such that the energy of this state, $\epsilon_{i}=\langle i|\hat{H}^{(\mu)}|i\rangle$, is closest to the mean energy, $\epsilon_{i}\approx E_{\rm av}^{(\mu)}$. We hence make the expression in Eq. (S8) more precise by explicitly referring to the expectation value of observable for a single Hamiltonian realization $H^{(\mu)}$,
$\mathcal{Q}(t)^{(\mu)}=\langle\psi_{0}|e^{iH^{(\mu)}t}\hat{S}_{L}^{z}e^{-iH^{(\mu)}t}|\psi_{0}\rangle\;,$ (S9)
setting $\hbar\equiv 1$. One can rewrite Eq. (S9) as
$\mathcal{Q}(t)^{(\mu)}=\sum_{n,m}c_{n}^{(\mu)*}c_{m}^{(\mu)}(S_{L}^{z})_{nm}e^{-i(E_{m}^{(\mu)}-E_{n}^{(\mu)})t}\;,$ (S10)
@@
$\mathcal{Q}_{\infty}^{(\mu)}\equiv\lim_{t\rightarrow\infty}\mathcal{Q}(t)^{(\mu)}=\sum_{n=1}^{\mathcal{D}}|c_{n}^{(\mu)}|^{2}(S_{L}^{z})_{n}\;.$ (S11)
@@
$\langle\mathcal{Q}^{(\mu)}\rangle_{\rm mic}\equiv\frac{1}{\mathcal{N}_{\epsilon_{i},\Delta_{\epsilon}}}\sum_{|E_{n}^{(\mu)}-\epsilon_{i}|<\Delta_{\epsilon}}(S_{L}^{z})_{n},$ (S12)
@@
where $\mathcal{N}_{\epsilon_{i},\Delta_{\epsilon}}$ is the normalization factor and the width $\Delta_{\epsilon}$ of the microcanonical ensemble is sufficiently small. In practice, we either set $\Delta_{\epsilon}=5\times 10^{-2}$, or, if the sum in Eq. (S12) contains too few elements (which may be the case for small systems), we select $\mathcal{N}_{\epsilon_{i},\Delta_{\epsilon}}=10$ eigenstates with energies $E_{n}^{(\mu)}$ closest to $\epsilon_{i}$, regardless of $\Delta_{\epsilon}$.
Figure S6: Scaling of the variance of temporal fluctuations $\sigma_{t}^{2}$, see Eq. (S16), vs $L$ for the observables (a) $\hat{S}^{z}_{L}$ and (b) $\hat{S}^{z}_{L-1}\hat{S}^{z}_{L}$. Results are shown for different values of $\alpha$.
Figure S7: Fluctuation exponents $\eta$ vs $\alpha$. Exponents of the fluctuations of the diagonal (squares) and the low-$\omega$ off-diagonal matrix elements (circles) are identical to those in Fig. 3(d). Exponent of the temporal fluctuations (triangles) are obtained by fitting the function $b_{0}2^{-2L/\eta}$ to the results in Fig. S6.
We define the averaged differences between the diagonal and the microcanonical ensembles as
$\Delta\mathcal{Q}_{\infty}={\rm Avr}_{\mu}\left\{|\mathcal{Q}_{\infty}^{(\mu)}-\langle\mathcal{Q}^{(\mu)}\rangle_{\rm mic}|\right\}\;,$ (S13)
where ${\rm Avr}_{\mu}\{...\}$ corresponds to the average over Hamiltonian realizations. We studied $\Delta\mathcal{Q}_{\infty}$ vs $1/L$ in the inset of Fig. 4 of the main text. The averaging over Hamiltonian realizations is carried out over $N_{\rm samples}=500$ for $L\leq 8$, and the number of realizations is then gradually reduced to $N_{\rm samples}=500-100(L-8)$ for $12>L\geq 9$. For $L=12,13$, we use $N_{\rm samples}=25,10$, respectively.
@@
We note that the initial value $\mathcal{Q}(t=0)^{(\mu)}$ at the quantum quench is either $+1/2$ or $-1/2$, depending on the polarization of the spin at $j=L$ in the initial state $|\psi_{0}\rangle$. To study properties of time evolution after averaging over different Hamiltonian realizations, we then define the autocorrelation function
$C(t)^{(\mu)}\equiv\langle\psi_{0}|\hat{S}_{L}^{z}(t)\hat{S}_{L}^{z}(0)|\psi_{0}\rangle\;,$ (S14)
@@
which is related to ${\cal Q}(t)^{(\mu)}$ from Eq. (S9) as $C(t)^{(\mu)}=s_{L}^{z}{\cal Q}(t)^{(\mu)}$, with $s_{L}^{z}=\langle\psi_{0}|\hat{S}_{L}^{z}|\psi_{0}\rangle$. In the main panel of Fig. 4 in the main text, we show the average autocorrelation function $C(t)$, which corresponds to $C(t)^{(\mu)}$ from Eq. (S14) averaged over different Hamiltonian realizations.
Another important property of quantum dynamics is whether or not the expectation values of observables equilibrate. To measure that, we define the variance of long-time temporal fluctuations for a given Hamiltonian realization dalessio_kafri_16,
$\sigma^{2}_{t,\mu}\equiv{\rm Var}_{t}\left\{\tilde{\mathcal{Q}}(t)^{(\mu)}\right\}\;,$ (S15)
where $\tilde{\mathcal{Q}}(t)^{(\mu)}$ is the expectation value $\mathcal{Q}(t)^{(\mu)}$ [Eq. (S14)] restricted to a time window after the Heisenberg time $t_{\rm H}$. Here, we select the time window $t\in[50t_{H},100t_{H}]$. The corresponding average variance over the Hamiltonian realizations is defined as
$\sigma^{2}_{t}={\rm Avr}_{\mu}\left\{\sigma^{2}_{t,\mu}\right\}\;.$ (S16)
One can set an upper bound for $\sigma^{2}_{t,\mu}$ that is given by the decay of the maximal off-diagonal matrix element with system size dalessio_kafri_16. This suggest a relationship between the temporal fluctuations and the fluctuations of matrix elements.
Here we ask whether the breakdown of the conventional ETH in the fading ergodicity regime is also manifested in the scaling of the variance of temporal fluctuations with system size. In Fig. S6 we show $\sigma_{t}^{2}$ vs $L$ for the observables $\hat{S}_{L}^{z}$ and $\hat{S}_{L-1}^{z}\hat{S}_{L}^{z}$, and different values of $\alpha$. We observe that in all cases under consideration, the decay of $\sigma_{t}^{2}$ is exponential with $L$, suggesting equilibration of the system. Moreover, we also observe that the rate of the exponential decay decreases when $\alpha$ approaches the ergodicity breaking transition, which is consistent with the softening of fluctuations of matrix elements. In Fig. S7 we compare the corresponding fluctuation exponents of the temporal fluctuations and the matrix elements fluctuations [as reported in Fig. 3(d)]. The obtained temporal fluctuation exponent is quantitatively smaller than the matrix element fluctuation exponent. However, in all cases, the fluctuation exponent increase when $\alpha$ approaches the transition point $\alpha_{c}$, suggesting that the long-time temporal fluctuations indeed carry information about the softening of the low-$\omega$ off-diagonal matrix elements.
Figure S8: Time evolution of the entanglement entropy ${\cal S}_{L}(t)$, see Eq. (S17), of a single (most weakly coupled) spin. Main panel: $\alpha=0.86>\alpha_{c}$ in the ergodic regime. Inset: $\alpha=0.72<\alpha_{c}$ in the non-ergodic regime. Results are shown for various system sizes $L$. Solid lines represent the averaged entanglement entropy ${\cal S}_{L}(t)$, while the temporal fluctuations in different Hamiltonian realizations are denoted by fading colors.
S6 Growth of entanglement entropy
To complement the analysis, we study the time evolution of the von Neumann entanglement entropy ${\cal S}_{L}(t)$, defined as
[…]
Results shown in Fig. S8 represent the entanglement growth in the ergodic regime (main panel) and in the non-ergodic regime (inset). We observe that in the ergodic regime, the entanglement entropy grows towards its maximum value, ${\cal S}(t\to\infty)\to\ln 2$. The approach towards maximal entanglement entropy is consistent with the ergodic behavior of the system (however, it is not a sufficient condition; see, e.g., Ref. vidmar_hackl_17 for a counter-example). On the other hand, the breakdown of ergodicity is clearly manifested in the inset of Fig. S8 at $\alpha<\alpha_{c}$, in which the entanglement entropy after a long time saturates to a value that is below the maximal value.

### v2

@@ Appendix D Distribution of matrix elements
Figure 9: PDFs $P(|O_{nn}|)$ of the diagonal matrix elements of observables: (a) $\hat{O}=\hat{S}^{z}_{L}$ and (b) $\hat{O}=\hat{S}^{z}_{\rm dot}\hat{S}^{z}_{L}$, for the system size $L=12$. Dashed lines indicate fits to the Gaussian PDF from Eq. (19). We numerically subtract the smooth structure function $O(\bar{E})$ to study the distribution of fluctuations.
Figure 10: Measures of gaussianity for the diagonal matrix elements of $\hat{S}_{L}^{z}$ vs $\alpha$, for different system sizes $L$. (a) The normalized variance $\Gamma_{O}-\pi/2$ from Eq. (20), (b) the Binder cumulant $B-1$ from Eq. (21), and (c) the Kurtosis $K-3$ from Eq. (22). Horizontal dotted lines are the results for the Gaussian PDF. Vertical dashed lines denote the transition point $\alpha_{c}$ obtained from Fig. 3(d) in the main text.
Figure 11: PDFs of off-diagonal matrix elements of the observable $\hat{O}=\hat{S}_{L}^{z}$, studied in the energy interval $\omega<100\Delta$, for (a) $\alpha=0.86$ and (b) $\alpha=0.92$. Main panels: $P(|O_{nm}|)$ vs $|O_{nm}|$. Dashed lines are the fits of the Gaussian PDF from Eq. (19) to the largest system size. Insets: $P(\ln|O_{nm}|)$ vs $\ln|O_{nm}|$. Solid lines are the fits of the corresponding Gaussian PDF from Eq. (23) using the same variance.
Figure 9 shows the distributions of the absolute values of diagonal matrix elements of observables $\hat{S}_{L}^{z}$ and $\hat{S}_{\rm dot}^{z}\hat{S}^{z}_{L}$, for a fixed system size and for different values of $\alpha$. Their main feature is that the distribution broadens upon decreasing $\alpha$, which is consistent with the softening of fluctuations in the regime of fading ergodicity.
The dashed lines in Fig. 9 represent the Gaussian probability density functions (PDFs). Given a positive random variable $z$, the corresponding Gaussian PDF is given by
$\bar{P}(z)=\frac{2}{\sqrt{2\pi\sigma^{2}}}\exp{-\frac{z^{2}}{2\sigma^{2}}}\;,$ (19)
@@
While the Gaussian PDFs in Fig. 9 provide accurate descriptions for the numerical distributions at large $\alpha$, they also exhibit deviations at lower $\alpha$ in the tails of the distributions. To quantify the deviations from the Gaussian PDF as functions of both $\alpha$ and $L$, we compute the normalized variance leblond_mallayya_19
$\Gamma_{O}=\frac{\expectationvalue{z^{2}}}{\expectationvalue{z}^{2}}\;,$ (20)
which yields $\Gamma_{O}=\pi/2$ for a Gaussian PDF from Eq. (19). We complement this measure by computing the Binder cumulant
$B=\frac{\expectationvalue{z^{4}}}{3\expectationvalue{z^{2}}^{2}}\;,$ (21)
@@
$K=\frac{\expectationvalue{(z-\expectationvalue{z})^{4}}}{\expectationvalue{(z-\expectationvalue{z})^{2}}^{2}}\;,$ (22)
@@
Figure 10 shows the dependence of the gaussianity measures from Eqs. (20)-(22) on $\alpha$, for the observable $\hat{S}_{L}^{z}$ and different system sizes $L$. We observe that the PDFs are indeed close to a Gaussian PDF at large $\alpha$, and notable deviations occur when the transition point is approached. However, the dependence on system size shares certain similarities with other ergodicity measures such as the average gap ratio $r$ studied in Fig. 7(b). This suggests a possibility that in the thermodynamic limit $L\to\infty$, the PDFs are close to Gaussian in the entire ergodic phase. The vertical dashed line in Fig. 10 represents the critical value $\alpha_{c}$ obtained by fitting the divergence of the fluctuation exponent $\eta$ in Fig. 3 using Eq. (10). Notably, all the gaussianity measures studied in Fig. 10 exhibit a scale invariant point that is very close to the extracted value of $\alpha_{c}$ from Fig. 3.
@@
$\tilde{P}(x)=\frac{2e^{x}}{\sqrt{2\pi\sigma^{2}}}\exp{-\frac{e^{2x}}{2\sigma^{2}}}\;.$ ) . (23)
The results from the insets of Fig. 11 reinforce the observation from the main panels that the deviations from the Gaussian PDF occur mostly in the tails, i.e., for large values of matrix elements. The deviations are larger in Fig. 11(a) [at $\alpha=0.86$] than in Fig. 11(b) [at $\alpha=0.92$], suggesting similar finite-size phenomenology as observed in Fig. 10 for the diagonal matrix elements.
Appendix E Details on quantum dynamics
In Sec. IV.2 of the main text, we studied the dynamics of the expectation value of observable $\hat{S}_{L}^{z}$ after a quantum quench, see Eq. (11). We choose the initial state $|\psi_{0}\rangle$ such that we minimize the finite-size effects due to the existence of a many-body mobility edge in the quantum sun model pawlik_sierant_2024 . Specifically, for a given Hamiltonian realization, denoted as $\hat{H}^{(\mu)}$, we select a single initial state $|\psi_{0}\rangle$ that is a product state in the computational basis $|i\rangle$ (i.e., an eigenstate of $\hat{S}_{L}^{z}$), such that the energy of this state, $\epsilon_{i}=\langle i|\hat{H}^{(\mu)}|i\rangle$, is closest to the mean energy, $\epsilon_{i}\approx E_{\rm av}^{(\mu)}$. Formally, Eq. (11) can be expressed as
$\mathcal{Q}(t)^{(\mu)}=\langle\psi_{0}|e^{iH^{(\mu)}t}\hat{S}_{L}^{z}e^{-iH^{(\mu)}t}|\psi_{0}\rangle\;,$ (24)
and further rewritten as
$\mathcal{Q}(t)^{(\mu)}=\sum_{n,m}c_{n}^{(\mu)*}c_{m}^{(\mu)}(S_{L}^{z})_{nm}e^{-i(E_{m}^{(\mu)}-E_{n}^{(\mu)})t}\;,$ (25)
@@ Appendix E Details on quantum dynamics
$\mathcal{Q}_{\rm DE}^{(\mu)}\equiv\lim_{t\rightarrow\infty}\mathcal{Q}(t)^{(\mu)}=\sum_{n=1}^{\mathcal{D}}|c_{n}^{(\mu)}|^{2}(S_{L}^{z})_{n}\;.$ (26)
@@
$\mathcal{Q}^{(\mu)}_{\rm ME}\equiv\frac{1}{\mathcal{N}_{\epsilon_{i},\Delta_{\epsilon}}}\sum_{|E_{n}^{(\mu)}-\epsilon_{i}|<\Delta_{\epsilon}}(S_{L}^{z})_{n},$ (27)
@@
where $\mathcal{N}_{\epsilon_{i},\Delta_{\epsilon}}$ is the normalization factor and the width $\Delta_{\epsilon}$ of the microcanonical ensemble is sufficiently small. In practice, we either set $\Delta_{\epsilon}=5\times 10^{-2}$, or, if the sum in Eq. (27) contains too few elements (which may be the case for small systems), we select $\mathcal{N}_{\epsilon_{i},\Delta_{\epsilon}}=10$ eigenstates with energies $E_{n}^{(\mu)}$ closest to $\epsilon_{i}$, regardless of $\Delta_{\epsilon}$. We then define the averaged differences between the diagonal and the microcanonical ensembles via Eq. (12), which is studied vs $L$ in Fig. 4(a) of the main text. The averaging over Hamiltonian realizations in Eq. (12) is carried out over $N_{\rm samples}=500$ for $L\leq 8$, and the number of realizations is then gradually reduced to $N_{\rm samples}=500-100(L-8)$ for $12>L\geq 9$. For $L=12,13$, we use $N_{\rm samples}=25,10$, respectively.
@@
We note that the initial value $\mathcal{Q}(t=0)^{(\mu)}$ at the quantum quench is either $+1/2$ or $-1/2$, depending on the polarization of the spin at $j=L$ in the initial state $|\psi_{0}\rangle$. To study properties of time evolution after averaging over different Hamiltonian realizations, we then introduced the spin-spin autocorrelation function $C(t)$ in Eq. (13). We calculate $C(t)$ by first defining the autocorrelation function for a single Hamiltonian realization,
$C(t)^{(\mu)}=\langle\psi_{0}|\hat{S}_{L}^{z}(t)\hat{S}_{L}^{z}(0)|\psi_{0}\rangle\;,$ (28)
@@
which is related to ${\cal Q}(t)^{(\mu)}$ from Eq. (11) as $C(t)^{(\mu)}=s_{L}^{z}{\cal Q}(t)^{(\mu)}$, with $s_{L}^{z}=\langle\psi_{0}|\hat{S}_{L}^{z}|\psi_{0}\rangle$. We then obtain $C(t)$ by averaging over different Hamiltonian realizations,
$C(t)={\rm Avr}_{\mu}\{C(t)^{(\mu)}\}\;.$ (29)
The time evolution of $C(t)$ is shown for $\alpha=0.86$ in Fig. 4(b) in the main text.